# Self-Tuning Spectral Clustering for Speaker Diarization

### v1

@@ Abstract
Spectral clustering has proven effective in grouping speech representations for speaker diarization tasks, although post-processing the affinity matrix remains difficult due to the need for careful tuning before constructing the Laplacian. In this study, we present a novel pruning algorithm to create a sparse affinity matrix called spectral clustering on p-neighborhood retained affinity matrix (SC-pNA). Our method improves on node-specific fixed neighbor selection by allowing a variable number of neighbors, eliminating the need for external tuning data as the pruning parameters are derived directly from the affinity matrix. SC-pNA does so by identifying two clusters in every row of the initial affinity matrix, and retains only the top $p\%$ similarity scores from the cluster containing larger similarities. Spectral clustering is performed subsequently, with the number of clusters determined as the maximum eigengap. Experimental results on the challenging DIHARD-III dataset highlight the superiority of SC-pNA, which is also computationally more efficient than existing auto-tuning approaches.
@@ I Introduction
The availability of precise annotations of audio recordings based on speaker information can significantly enhance the field of audio analytics involving multi-speaker conversations. This has applications across various domains, from everyday online meetings to advanced speech and audio forensics. This task is formally known as speaker diarization (SD), where the goal is to assign speaker labels to segments of speech based on speaker identity [1]. Traditionally, SD addresses the problem of “who spoke when” in a given audio recording, particularly in multi-speaker environments. SD pipeline conventionally consists of several interconnected components, including speech enhancement (SE), speech activity detection (SAD), segmentation, speaker embedding extraction, clustering, and re-segmentation.
@@
Modern SD architectures based on deep learning [2] have demonstrated remarkable performance, surpassing classical methods on widely used benchmark datasets such as CALLHOME [3], AMI [4], and VoxSRC [5]. However, SD systems still face practical challenges when applied to more realistic conversational speech datasets, such as subsets of the DIHARD dataset [6]. Factors like background noise, overlapping speech, a large number of speakers, imbalanced speech contributions across speakers, intra-speaker variability, and the recording environment’s lack of domain adaptation make the task even more complex. While issues such as environmental noise, SAD, speech overlap, and intra-speaker variability are extensively studied, the clustering problem does not receive the necessary attention despite playing a crucial role in the overall diarization process.
@@
Clustering is a classical problem in unsupervised machine learning with a substantial body of literature [7, 8]. However, in the context of speaker diarization (SD), clustering algorithms are largely limited to basic techniques such as k-means, agglomerative hierarchical clustering, mean-shift [9], and spectral clustering (SC) [10]. Among these, SC is the most widely used, particularly in modern SD systems based on deep speaker embeddings [11]. This preference is primarily due to its simplicity, with fewer parameters, and its strong mathematical foundations [12, 13].
@@
The spectral clustering approach used for the SD task involves specific parameters during the adjustment of the affinity matrix. Since the nodes represented by speaker embeddings capture speaker characteristics, it is crucial to minimize the impact of different speaker similarities and to retain or enhance same-speaker similarities [11]. In practice, a sparse graph is formed by removing less similar nodes and setting their values to zero. This process requires a pruning parameter typically tuned using externally labeled speech data. However, audio from different domains may introduce domain mismatches, significantly degrading SD performance. Recently, an auto-tuning approach for SC, named as auto-tuning spectral clustering (ASC), has been proposed in the context of SD, eliminating the need for tuning on external data [11]. The tuning parameters are instead derived directly from the audio recording being evaluated, using a brute-force search across different parameter values and optimizing a proxy evaluation metric. While this technique is promising, it has several limitations. First, it requires repeated spectral clustering computations involving eigenvalue decomposition, making it computationally expensive for longer recordings. Second, the proxy evaluation metric has limitations and does not always accurately reflect actual diarization performance [11]. Third, it uses a fixed number of neighbors for each node, which is problematic, especially for audio recordings with an unbalanced distribution of speech. Additionally, the graph may become disconnected in specific situations, leading to the failure of the spectral clustering algorithm or suboptimal results [13, 14].
@@
This work introduces a novel self-tuning method called Spectral clustering on p-neighborhood retained affinity matrix (SC-pNA) for speaker diarization, addressing the limitations of previous state-of-the-art (SOTA) adaptive spectral clustering approaches. Inspired by the score distribution characteristics reported in speaker recognition literature [15, 16], we propose a score-based, node-specific pruning strategy. In this method, the pruning threshold is dynamically adjusted for each node, allowing the retention of nodes with stronger connections while eliminating weakly connected, unreliable nodes. Like ASC, our approach does not require external data for threshold adjustment. However, it advances further by eliminating the need for proxy evaluation metrics. Additionally, it requires only a single eigenvalue decomposition of the Laplacian matrix, which is computed once from the refined adjacency matrix.
@@ II-A Conventional spectral clustering (CSC)
Spectral clustering  [13] is a technique used to group data points into clusters based on the eigenvalues and eigenvectors of a similarity matrix constructed from the data. It leverages the properties of the graph Laplacian to partition the data in a way that often captures complex cluster structures better than traditional methods like k-means. It has considerable applications in the area of image segmentation [14]. With the emergence of speech embeddings [17, 18, 19], SC is widely adopted for SD [10, 20, 21, 2]. The SC framework for SD is summarized as follows:
Given a finite set $\mathbf{X}=$$x_{1},x_{2},...,x_{N}$} of $N$ embeddings extracted from their corresponding $N$ speech segments. Our objective is to cluster them into $k$ speaker classes. First, an affinity matrix $\mathbf{M}\in\mathbb{R}^{N\times N}$ is computed. Each element in $\mathbf{M}$ represent the cosine similarity between speech embeddings. Unreliable similarity scores are pruned in each row of $\mathbf{M}$. A pruning parameter $\alpha$ determines the fixed $\left\lfloor N(1-\alpha)\right\rfloor$ entries to be made zero. The resulting matrix $\mathbf{M}_{\alpha}$ is symmetrized. Further, an unnormalized Laplacian matrix is computed as $\mathbf{L}=\mathbf{D}-\mathbf{M}_{\alpha}$, where $\mathbf{D}$ is a degree matrix. The standard $k$-means algorithm is used to cluster the first $k$ eigenvectors corresponding to higher eigenvalues obtained from the eigendecomposition of $\mathbf{L}$. The optimal pruning parameter $\alpha$ is estimated by tuning it on the development set. It is done by varying $\alpha$ linearly from 0 to 1 with a typical step size of 0.01. The value of $\alpha$ that exhibits the minimum diarization error rate (DER) is utilized for the evaluation set. But this setting is not the best choice, as in scenarios such as where the development and evaluation data can come from situations that are substantially different [22].
@@ II-B Auto-tuning spectral clustering (ASC)
To mitigate this data-dependency, another variant of SC known as auto-tuning spectral clustering (ASC) was proposed in [11]. This approach shares a fundamental similarity with CSC, as both require repeated computation of spectral clustering (SC) at varying pruning rates. However, rather than relying on actual DER measured with the ground truth from a labeled development set, this method estimates a proxy DER directly from the unlabelled data being diarized. The proxy DER is calculated by first determining the maximum eigengap, normalized by the largest eigenvalue, and then computing the ratio of the pruning factor to the maximum eigengap.
This technique is preferred over the CSC as it does not require any ground truth but it also has some limitations. It is computationally expensive, as it requires eigen-decomposition of the Laplacian matrix $\mathbf{L}$ as many times as the number of embeddings $N$. So, it is not a realistic choice in scenarios where the recordings are arbitrarily longer. Also, it works on the concept of a proxy-DER, which may always not be a good approximation of the original DER. Also, it prunes a fixed number of elements in each row of the affinity matrix, which is also not an ideal choice in situations when the distribution of spoken duration of speakers is not uniform across a recording.
@@ II-C Spectral clustering on $p$-neighborhood retained affinity matrix (SC-pNA)
Background: Our proposed method of SC-pNA is focused on creating a sparse affinity matrix that can aid the spectral clustering method to better delineate different speaker embeddings, thereby improving speaker diarization. To create this sparse affinity matrix, each row is pruned independently by removing low similarity scores. Each row of the original affinity matrix should ideally have high similarities to the embeddings of the same speakers, while having low similarities to the embeddings of all other speakers.
As the number of embeddings belonging to the same speaker will vary across a recording, therefore the number of similarity scores retained in each row should be decided adaptively. In general, the similarity scores in each row of an affinity matrix can be described in terms of two distributions. The first within-cluster distribution $C_{w}$ corresponds to the high similarity scores between same speaker embeddings, whereas the second between-cluster distribution $C_{b}$ describes the low similarity scores between different speaker embeddings. This approach of describing similarities in terms of two distributions has been previously explored in the context of speaker recognition [15, 23], where $C_{w}$ and $C_{b}$ are assumed to follow the Gaussian distribution, i.e., $C_{w}\sim\mathcal{N}(\mu_{w},\sigma_{w})$, and $C_{b}\sim\mathcal{N}(\mu_{b},\sigma_{b})$. As we wish to accept $C_{w}$ and reject $C_{b}$, the equal error rate (EER) provides a location where the false rejection rate of $C_{w}$ equals the false acceptance rate of $C_{b}$. The analytical EER is defined as,
@@
The theoretical foundations of the threshold $\Delta$ make it an appealing choice to delineate distributions $C_{w}$ and $C_{b}$ in each row of the affinity matrix, and thereafter all entries corresponding to $C_{b}$ can be set to zero to create the sparse affinity matrix. We refer to this method as the EER-$\Delta$ approach: on an initial affinity matrix $A$, $\Delta_{i}$ is first computed over each row $A_{i}$, for $i=1,...,n$. A sparse affinity matrix $B$ is then created by using $\Delta_{i}$ as a threshold in each row $B_{i}$, where we set $B_{ij}=0$ if $A_{ij}<\Delta_{i}$, otherwise $B_{ij}=A_{ij}$. Finally, spectral clustering is conducted on $B$, using the maximum eigengap to determine the number of clusters. The EER-$\Delta$ approach has a lower computation cost compared to the CSC and ASC as it does not require tuning. An additional advantage of EER-$\Delta$ is that the number of similarity scores pruned in each row can vary, depending on the sizes of the identified $C_{w}$ and $C_{b}$.
@@
SC-pNA Methodology: In SC-pNA, we consider an even more aggressive thresholding approach, to create a sparser affinity matrix. As the $\Delta_{i}$ thresholds of EER-$\Delta$ retain all similarity scores for $C_{w}$ in the $i$-th row, we can interpret this approach as restricting the neighborhood of the $i$-th speaker segment $x_{i}$ only to the $C_{w}$ cluster. Let $N(x_{i})$ be the set of speaker segments in the neighborhood of $x_{i}$, and let $C_{w}^{(i)}$ be the cluster identified for $x_{i}$. Then each $\Delta_{i}$ threshold of EER-$\Delta$ establishes,
$N(x_{i})=C_{w}^{(i)}\text{ for }i=1,2,...,n,$ (3)
@@
In the proposed SC-pNA, a more aggressive pruning is pursued, where only the top $p\%$ similarities from the identified $C_{w}^{(i)}\backslash\{x_{i}\}$ are retained in every row. This is shown in Figure 1, where the proposed pruning approach can be interpreted as considering higher false rejection ratio (FAR), with the thresholds empirically set at (1-p)FAR. The constant high self-similarity scores $s(x_{i},x_{i})=1$ are excluded from the process to identify the clusters $C_{w}^{(i)}$ and $C_{b}^{(i)}$, as they can behave as outliers when all other similarities $s(x_{i},x_{j}),\forall j\neq i$, are much smaller than $s(x_{i},x_{i})$. This exclusion is interpreted as the removal of all self-loops, thus leading to an adjacency matrix where connectivity is maintained only through neighborhood connections. Thus retaining only top $p\%$ similarities creates the neighborhoods for all $x_{i}$,
$N_{p}(x_{i})=S^{(i)}\subseteq C_{w}^{(i)}\backslash\{x_{i}\}.$ (4)
Comparing equations (3) and (4) lets us determine that $|N_{p}(x_{i})|\leq|N(x_{i})|$, and thus smaller neighborhoods are formed, leading to more sparse affinity matrices. However, these smaller neighborhoods should be capable of identifying the correct speaker clusters. If a recording contains $k$ speakers, the embeddings should be clustered correctly into $k$ clusters $C_{1},\ldots,C_{k}$, where each cluster $C_{j}=\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$ contains $|C_{j}|$ embeddings. Thus $p$ should be chosen so that $|N_{p}(x_{i})|$ is small, yet cluster $C_{j}$ can be recovered by examining the neighborhoods of $\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$, i.e.,
$N_{p}(x_{j1})\cup N_{p}(x_{j2})\cup\ldots\cup N_{p}(x_{j|C_{j}|})=C_{j}.$ (5)
In the EER-$\Delta$ approach, $N(x_{j1})\cup\ldots\cup N(x_{j|C_{j}|})=C_{j}$ holds trivially. For real world data, that can be noisy and have overlapped clusters, the challenge for both SC-pNA and the EER-$\Delta$ approach lies in utilising the available neighborhood information to identify the set of embeddings $\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$ that recovers each cluster $C_{j}$ accurately.
@@
    Create the adjacency matrix without self-loops: The initial adjacency matrix $A$ is formed using cosine similarity,
    $A_{ij}=\frac{x_{i}^{T}x_{j}}{||x_{i}||\,||x_{j}||},\ \forall i,j=1,2,\ldots,n.$ (6)
    The diagonal entries of $A$ are set to zero i.e., $A_{ii}=0,\ i=1,2,\ldots,n$.
@@
    Identify $C_{w}^{(i)}\ \text{and}\ C_{b}^{(i)}$ in each row: In each row $A_{i}$, two clusters $C_{1}^{(i)},C_{2}^{(i)}$ are identified using $k$-Means clustering with $k=2$. From these two clusters, $C_{w}^{(i)}$ is identified as the one with the larger cluster center,
@@
    Retain top $p\%$ similarity scores: A matrix $P$ is formed which retains in each row the top $p\%$ similarity scores from the cluster $C_{w}^{(i)}$, and the rest are set to zero, i.e.,
    $P_{ij}=\begin{cases}0&\text{, if }A_{ij}\text{ not in the top }p\%\text{ of }C_{w}^{(i)}\\ A_{ij}&\text{, otherwise}\end{cases}.$ (8)
@@
    Symmetrize and form the Laplacian: A symmetric matrix $W$ is formed from $P$,
    $W=\frac{1}{2}(P+P^{T}).$ (9)
    A diagonal matrix $D$ is created with the sum of the rows of $W$ as its entries: $D_{ii}=\sum_{j=1}^{n}|w_{ij}|$. The Laplacian $L$ is then created,
    $L=D-W.$ (10)
@@
    Compute eigengap and estimate $\hat{k}$ : For each recording, a user can specify a reasonable assumption for the maximum number of speakers $k_{\max}$. We find the smallest $M$ eigenvalues $\lambda_{1},\ldots,\lambda_{M}$ of $L$, where $M=\min\{k_{\max},n\}$. The $M-1$ length eigengap vector is constructed as,
@@
    Cluster the spectral embeddings: Form matrix $V\in\mathbb{R}^{n\times\hat{k}}$, containing in its columns the $\hat{k}$ eigenvectors corresponding to the smallest $\hat{k}$ eigenvalues of $L$. Apply, $k$-Means clustering on the rows of $V$ to obtain cluster memberships $\hat{y}\in[1,2,\ldots,\hat{k}]^{n}$.
The algorithm for the proposed SC-pNA is outlined in Algorithm 1. As SC-pNA does not require any tuning, and only performs a single eigenvalue decomposition of the Laplacian matrix, it incurs significantly lower computation costs compared to the SOTA ASC approach.
@@
  Output: Cluster memberships $\hat{y}\in[1,2,\ldots,\hat{k}]^{n}$.
  1. Compute $A$ using equation (6).
@@
  6. Form $P$ using equation (8).
  7. Form $W$ using equation (9).
  8. Form $L$ using equation (10).
@@
  11. Obtain $\hat{y}$ using $k$-Means on the spectral embeddings in $V$, corresponding to the $\hat{k}$ smallest eigenvalues of $L$.

### v2

@@ Abstract
Spectral clustering has proven effective in grouping speech representations for speaker diarization tasks, although post-processing the affinity matrix remains difficult due to the need for careful tuning before constructing the Laplacian. In this study, we present a novel pruning algorithm to create a sparse affinity matrix called spectral clustering on p-neighborhood retained affinity matrix (SC-pNA). Our method improves on node-specific fixed neighbor selection by allowing a variable number of neighbors, eliminating the need for external tuning data as the pruning parameters are derived directly from the affinity matrix. SC-pNA does so by identifying two clusters in every row of the initial affinity matrix, and retains only the top $p\%$ similarity scores from the cluster containing larger similarities. Spectral clustering is performed subsequently, with the number of clusters determined as the maximum eigengap. Experimental results on the challenging DIHARD-III dataset highlight the superiority of SC-pNA, which is also computationally more efficient than existing auto-tuning approaches. Our implementations are available at https://github.com/nikhilraghav29/SC-pNA.
@@ I Introduction
The availability of precise annotations of audio recordings based on speaker information can significantly enhance the field of audio analytics involving multi-speaker conversations. This has applications across various domains, from everyday online meetings to advanced speech and audio forensics. This task is formally known as speaker diarization (SD), where the goal is to assign speaker labels to segments of speech based on speaker identity [1]. Traditionally, SD addresses the problem of “who spoke when” in a given audio recording, particularly in multi-speaker environments. The SD pipeline conventionally consists of several interconnected components, including speech enhancement (SE), speech activity detection (SAD), segmentation, speaker embedding extraction, clustering, and re-segmentation.
@@
Modern SD architectures based on deep learning [2] have demonstrated remarkable performance, surpassing classical methods on widely used benchmark datasets such as CALLHOME [3], AMI [4], and VoxSRC [5]. However, SD systems still face practical challenges when applied to more realistic conversational speech datasets, such as the DIHARD-III dataset [6]. Factors like background noise, overlapping speech, a large number of speakers, imbalanced speech contributions across speakers, intra-speaker variability, and variability in the recording environments make the task even more complex. While issues such as environmental noise, SAD, speech overlap, and intra-speaker variability are extensively studied, the clustering problem does not receive the necessary attention despite playing a crucial role in the overall diarization process.
@@
Clustering is a classical problem in unsupervised machine learning with a substantial body of literature [7, 8]. However, in the context of SD, clustering algorithms are largely limited to basic techniques such as k-means, agglomerative hierarchical clustering, mean-shift [9], and spectral clustering (SC) [10]. Among these, SC is the most widely used, particularly in modern SD systems based on deep speaker embeddings [11]. This preference is primarily due to its simplicity, with fewer parameters, and its strong mathematical foundations [12, 13].
@@
The spectral clustering approach used for the SD task involves specific parameters during the adjustment of the affinity matrix. Since the nodes represented by speaker embeddings capture speaker characteristics, it is crucial to minimize the impact of different speaker similarities and to retain or enhance same-speaker similarities [11]. In practice, a sparse graph is formed by removing less similar nodes and setting their values to zero. This process requires a pruning parameter typically tuned using a development set containing labeled speech data. However, audio from different domains may introduce domain mismatches, substantially degrading SD performance. Recently, an auto-tuning approach for SC, named as auto-tuning spectral clustering (ASC), has been proposed in the context of SD, eliminating the need for tuning on external data [11]. The tuning parameters are instead derived directly from the audio recording being evaluated, using a brute-force search across different parameter values and optimizing a proxy evaluation metric. While this technique is promising, it has several limitations. First, it requires repeated spectral clustering computations involving eigenvalue decomposition, making it computationally expensive for longer recordings. Second, the proxy evaluation metric has limitations and does not always accurately reflect actual diarization performance [11]. Third, it uses a fixed number of neighbors for each node, which is problematic, especially for audio recordings with an imbalanced distribution of speech. Additionally, the graph may become disconnected in specific situations, leading to the failure of the spectral clustering algorithm, or may otherwise produce suboptimal results [13, 14].
@@
This work introduces a novel self-tuning method called spectral clustering on p-neighborhood retained affinity matrix (SC-pNA) for SD, addressing the limitations of previous state-of-the-art (SOTA) adaptive spectral clustering approaches. Inspired by the score distribution characteristics reported in speaker recognition literature [15, 16], we propose a score-based, node-specific pruning strategy. In this method, the pruning threshold is dynamically adjusted for each node, allowing the retention of nodes with stronger connections while eliminating weakly connected, unreliable nodes. Like ASC, our approach does not require external data for threshold adjustment. However, it advances further by eliminating the need for proxy evaluation metrics. Additionally, it requires only a single eigenvalue decomposition of the Laplacian matrix, which is computed once from the refined adjacency matrix.
@@ II-A Conventional spectral clustering (CSC)
Spectral clustering  [13, 14] is a technique used to group data points into clusters based on the eigenvalues and eigenvectors of a similarity matrix constructed from the data. It leverages the properties of the graph Laplacian to partition the data in a way that often captures complex cluster structures better than traditional methods like $k$-means. With the emergence of speech embeddings [17, 18, 19], SC is widely adopted for SD [10, 20, 21, 2].
The semi-supervised CSC [2] for SD aims to identify $k$ speaker classes from a finite set $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$ of $n$ embeddings, extracted from $n$ speech segments. CSC first computes an affinity matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$, containing the pairwise cosine similarities between speech embeddings. Next, $\left\lfloor n(1-\alpha)\right\rfloor$ similarity scores are deemed unreliable and hence pruned in every row of $\mathbf{M}$, using a single pruning parameter $\alpha$. The resulting matrix $\mathbf{M}_{\alpha}$ is symmetrized, and an unnormalized Laplacian matrix is computed as $\mathbf{L}=\mathbf{D}-\mathbf{M}_{\alpha}$, where $\mathbf{D}$ is a degree matrix. The number of speakers $k$ is estimated using the eigengap approach [13], and finally, the $k$-means algorithm clusters the $k$ eigenvectors of $\mathbf{L}$ corresponding to its $k$ smallest eigenvalues. CSC is semi-supervised, as the optimal pruning parameter $\alpha$ exhibiting minimum diarization error rate (DER) is estimated from a development set, by varying $\alpha$ linearly from 0 to 1 with a typical step size of 0.01. This setting is however not the best choice when the development and evaluation data comes from substantially different environments [22].
@@ II-B Auto-tuning spectral clustering (ASC)
To not depend on a development set for parameter tuning, an unsupervised variant of SC known as auto-tuning spectral clustering (ASC) was proposed in [11]. ASC estimates a proxy DER from the unlabelled data. This is done by first determining the maximum eigengap, normalized by the largest eigenvalue, and then computing the ratio of the pruning factor to the maximum eigengap. The proxy DER however may not always be a good approximation of the original DER. Additional limitations of both ASC and CSC are their high computation costs. This is due to the repeated eigendecomposition of the Laplacian formed at varying pruning rates, which is required to estimate the maximum eigengap. Both methods also prune a strictly fixed number of entries across all rows of the affinity matrix, which is also not an ideal choice in situations when the distribution of spoken duration of speakers is not uniform across a recording.
@@ II-C Spectral clustering on $p$-neighborhood retained affinity matrix (SC-pNA)
Background: Our proposed method of SC-pNA is focused on creating a sparse affinity matrix that can aid the spectral clustering method to better delineate different speaker embeddings, thereby improving SD. To create this sparse affinity matrix, each row is pruned independently by removing low similarity scores. Each row of the original affinity matrix should ideally have high similarities to the embeddings of the same speakers, while having low similarities to the embeddings of all other speakers.
As the number of embeddings belonging to the same speaker will vary across a recording, therefore the number of similarity scores retained in each row should be decided adaptively. In general, the similarity scores in each row of an affinity matrix can be described in terms of two distributions. The first within-cluster distribution $C_{w}$ corresponds to the high similarity scores between the same speaker embeddings, whereas the second between-cluster distribution $C_{b}$ describes the low similarity scores between different speaker embeddings. This approach of describing similarities in terms of two distributions has been previously explored in the context of speaker recognition [15, 23], where $C_{w}$ and $C_{b}$ are assumed to follow the Gaussian distribution, i.e., $C_{w}\sim\mathcal{N}(\mu_{w},\sigma_{w})$, and $C_{b}\sim\mathcal{N}(\mu_{b},\sigma_{b})$. As we wish to accept $C_{w}$ and reject $C_{b}$, the equal error rate (EER) provides a location where the false rejection rate of $C_{w}$ equals the false acceptance rate of $C_{b}$. The analytical EER is defined as,
@@
The theoretical foundations of the threshold $\Delta$ make it an appealing choice to delineate distributions $C_{w}$ and $C_{b}$ in each row of the affinity matrix, and thereafter all entries corresponding to $C_{b}$ can be set to zero to create the sparse affinity matrix. We refer to this method as the EER-$\Delta$ approach: on an initial affinity matrix $\mathbf{A}$, $\Delta_{i}$ is first computed over each row $\mathbf{A}_{i}$, for $i=1,...,n$. A sparse affinity matrix $\mathbf{B}$ is then created by using $\Delta_{i}$ as a threshold in each row $\mathbf{B}_{i}$, where we set $\mathbf{B}_{ij}=0$ if $\mathbf{A}_{ij}<\Delta_{i}$, otherwise $\mathbf{B}_{ij}=\mathbf{A}_{ij}$. Finally, spectral clustering is conducted on $\mathbf{B}$, using the maximum eigengap to determine the number of clusters. The EER-$\Delta$ approach has a lower computation cost compared to the CSC and ASC, as it only computes the eigendecomposition of a single Laplacian, formed from the affinity matrix whose rows are pruned using $\Delta_{i}$. An additional advantage of EER-$\Delta$ is that the number of similarity scores pruned in each row is adaptive, and depends on the sizes of the identified $C_{w}$ and $C_{b}$.
@@
SC-pNA Methodology: In SC-pNA, we consider an even more aggressive thresholding approach, to create a sparser affinity matrix. As the $\Delta_{i}$ thresholds of EER-$\Delta$ retain all similarity scores for $C_{w}$ in the $i$-th row, we can interpret this approach as restricting the neighborhood of the $i$-th speaker segment $\mathbf{x}_{i}$ only to the $C_{w}$ cluster. Let $N(\mathbf{x}_{i})$ be the set of speaker segments in the neighborhood of $\mathbf{x}_{i}$, and let $C_{w}^{(i)}$ be the cluster identified for $\mathbf{x}_{i}$. Then each $\Delta_{i}$ threshold of EER-$\Delta$ establishes,
$N(\mathbf{x}_{i})=C_{w}^{(i)}\text{ for }i=1,2,...,n,$ (3)
@@
In the proposed SC-pNA, a more aggressive pruning is pursued, where only the top $p\%$ similarities from the identified $C_{w}^{(i)}\backslash\{\mathbf{x}_{i}\}$ are retained in every row. This is shown in Figure 1, where the proposed pruning approach can be interpreted as considering higher false rejection ratio (FRR), with the thresholds empirically set at (1-p)FRR. The constant high self-similarity scores $s(\mathbf{x}_{i},\mathbf{x}_{i})=1$ are excluded from the process to identify the clusters $C_{w}^{(i)}$ and $C_{b}^{(i)}$, as they can behave as outliers when all other similarities $s(\mathbf{x}_{i},\mathbf{x}_{j}),\forall j\neq i$, are much smaller than $s(\mathbf{x}_{i},\mathbf{x}_{i})$. This exclusion is interpreted as the removal of all self-loops, thus leading to an adjacency matrix where connectivity is maintained only through neighborhood connections. Thus retaining only top $p\%$ similarities creates the neighborhoods for all $\mathbf{x}_{i}$,
$N_{p}(\mathbf{x}_{i})\subseteq C_{w}^{(i)}\backslash\{\mathbf{x}_{i}\}.$ (4)
Comparing equations (3) and (4) lets us determine that $|N_{p}(\mathbf{x}_{i})|\leq|N(\mathbf{x}_{i})|$, and thus smaller neighborhoods are formed, leading to more sparse affinity matrices. However, these smaller neighborhoods should be capable of identifying the correct speaker clusters. If a recording contains $k$ speakers, the embeddings should be clustered correctly into $k$ clusters $C_{1},\ldots,C_{k}$, where each cluster $C_{j}=\{\mathbf{x}_{j1},\mathbf{x}_{j2},\ldots,\mathbf{x}_{j|C_{j}|}\}$ contains $|C_{j}|$ embeddings. Thus $p$ should be chosen so that $|N_{p}(\mathbf{x}_{i})|$ is small, yet cluster $C_{j}$ can be recovered by examining the neighborhoods of $\{\mathbf{x}_{j1},\mathbf{x}_{j2},\ldots,\mathbf{x}_{j|C_{j}|}\}$, i.e.,
$N_{p}(\mathbf{x}_{j1})\cup N_{p}(\mathbf{x}_{j2})\cup\ldots\cup N_{p}(\mathbf{x}_{j|C_{j}|})=C_{j}.$ (5)
In the EER-$\Delta$ approach, $N(\mathbf{x}_{j1})\cup\ldots\cup N(\mathbf{x}_{j|C_{j}|})=C_{j}$ holds trivially. For real world data, that can be noisy and have overlapped clusters, the challenge for both SC-pNA and the EER-$\Delta$ approach lies in utilising the available neighborhood information to identify the set of embeddings $\{\mathbf{x}_{j1},\mathbf{x}_{j2},\dots,\mathbf{x}_{j|C_{j}|}\}$ that recovers each cluster $C_{j}$ accurately.
@@
    Create the adjacency matrix without self-loops: The initial adjacency matrix $\mathbf{A}$ is formed using cosine similarity,
    $\mathbf{A}_{ij}=\frac{\mathbf{x}_{i}^{\top}\mathbf{x}_{j}}{\|\mathbf{x}_{i}\|\|\mathbf{x}_{j}\|},\quad\forall i,j\in\{1,2,\ldots,n\},\ i\neq j.$ (6)
    The diagonal entries of $\mathbf{A}$ are set to zero i.e., $\mathbf{A}_{ii}=0,\ i=1,2,\ldots,n$.
@@
    Identify $C_{w}^{(i)}\ \text{and}\ C_{b}^{(i)}$ in each row: In each row $\mathbf{A}_{i}$, two clusters $C_{1}^{(i)},C_{2}^{(i)}$ are identified using $k$-Means clustering with $k=2$. From these two clusters, $C_{w}^{(i)}$ is identified as the one with the larger cluster center,
@@
    Retain top $p\%$ similarity scores: A matrix $\mathbf{P}$ is formed which retains in each row the top $p\%$ similarity scores from the cluster $C_{w}^{(i)}$, and the rest are set to zero, i.e.,
    $P_{ij}=\begin{cases}0&\text{, if }\mathbf{A}_{ij}\text{ is not in the top }p\%\text{ of }C_{w}^{(i)}\\ \mathbf{A}_{ij}&\text{, otherwise}\end{cases}.$ (8)
@@
    Symmetrize and form the Laplacian: A symmetric matrix $\mathbf{W}$ is formed from $\mathbf{P}$,
    $\mathbf{W}=\frac{1}{2}\left(\mathbf{P}+\mathbf{P}^{\top}\right).$ (9)
    A diagonal matrix $\mathbf{D}$ is created with the sum of the rows of $\mathbf{W}$ as its entries: $\mathbf{D}_{ii}=\sum_{j=1}^{n}|w_{ij}|$. The Laplacian $\mathbf{L}$ is then created,
    $\mathbf{L}=\mathbf{D}-\mathbf{W}.$ (10)
@@
    Compute eigengap and estimate $\hat{k}$ : For each recording, a user can specify a reasonable assumption for the maximum number of speakers $k_{\max}$. We find the smallest $M$ eigenvalues $\lambda_{1},\ldots,\lambda_{M}$ of $\mathbf{L}$, where $M=\min\{k_{\max},n\}$. The $M-1$ length eigengap vector is constructed as,
@@
    Cluster the spectral embeddings: Form matrix $\mathbf{V}\in\mathbb{R}^{n\times\hat{k}}$, containing in its columns the $\hat{k}$ eigenvectors corresponding to the smallest $\hat{k}$ eigenvalues of $\mathbf{L}$. Apply $k$-Means clustering on the rows of $\mathbf{V}$ to obtain cluster memberships $\hat{\mathbf{y}}\in[1,2,\ldots,\hat{k}]^{n}$.
Our proposed SC-pNA approach is outlined in Algorithm 1. SC-pNA does not require any tuning, and has an $O(n^{3})$ complexity due to step 9, while the remaining steps have $O(n^{2})$ cost or lower. In contrast, ASC has a higher complexity of $O(Pn^{3})=O(n^{4})$, where $P=\lfloor n/4\rfloor$ is the recommended number of binarized similarity matrices.
@@
  Output: Cluster memberships $\hat{\mathbf{y}}\in[1,2,\ldots,\hat{k}]^{n}$.
  1. Compute $\mathbf{A}$ using equation (6).
@@
  6. Form $\mathbf{P}$ using equation (8).
  7. Form $\mathbf{W}$ using equation (9).
  8. Form $\mathbf{L}$ using equation (10).
@@
  11. Obtain $\hat{\mathbf{y}}$ using $k$-Means on the spectral embeddings in $\mathbf{V}$, corresponding to the $\hat{k}$ smallest eigenvalues of $\mathbf{L}$.